# Are Biological Systems More Intelligent Than Artificial Intelligence?

### v6

@@ Abstract
Are biological self-organising systems more “intelligent” than artificial intelligence (AI)? If so, why? I address this question using a mathematical framework that defines intelligence in terms of adaptability. Systems are modelled as stacks of abstraction layers (Stack Theory) and compared by how effectively they delegate agentic control down their stacks. I illustrate this using computational, biological, military, governmental, and economic systems. Contemporary AI typically relies on static, human-engineered stacks whose lower layers are fixed during deployment. Put provocatively, such systems resemble inflexible bureaucracies that adapt only top-down. Biological systems are more intelligent because they delegate adaptation. Formally, I prove a theorem (The Law of the Stack) showing that adaptability at higher layers is bottlenecked by adaptability at lower layers. I further show that, under standard viability assumptions, maximising adaptability is equivalent to minimising variational free energy, implying that delegation is necessary for free-energy minimisation. Generalising bioelectric accounts of cancer as isolation from collective informational structures, I analyse cancer-like failure modes in non-biological systems when delegation is inadequate. This yields design principles for building robust systems via delegated control, and reframes hybrid agents (e.g. organoids or human–AI systems) as weak boundary-condition design problems in which constraints shape low-level policy spaces while preserving collective identity.
@@ 1 Introduction
Organoids are living systems whose behaviour is sculpted by engineered constraints. Human brain cells embedded in a computational interface have been trained to play Pong, achieving remarkable sample efficiency compared to a contemporary reinforcement learning agent [53]. Similarly, Physarum slime mould is normally an indiscriminate expander. However it reliably computes shortest paths when placed inside a maze [62]. These behaviours are neither programmed nor accidental. They arise because biological agency is coupled to carefully designed boundary conditions. Such systems are hybrid agents, jointly produced by bottom-up cellular dynamics and top-down environmental structure [82].
@@
Biological systems self-organise across scales. Their parts and environment interact to produce spatiotemporal order [4, 78, 76, 9, 51, 39, 92]. Many systems proactively learn, storing information about the past to anticipate future needs [54]. To adapt in complex environments, a system must infer cause and effect [72, 60, 7]. Throughout, I use ‘intelligence’ in the specific sense of the sample and energy efficiency of adaptation [16, 26, 11, 14] [^1] Mainstream AI typically treats intelligence as performance across tasks. My conclusions concern efficiency of acquiring said performance. Accordingly, claims about “AI” should be read as applying to dominant contemporary ML training and deployment stacks, rather than all possible systems. See ESM Note S1..
@@
Two common explanations for biological efficiency are active inference under the Free Energy Principle (FEP) [39, 40, 71] and powerful inductive biases [26]. These address, respectively, how agents update beliefs and act, and what structure is pre-specified. This paper isolates a complementary structural question, namely where in a system is adaptation permitted?
@@
Computers are organised into layers of abstraction. In dominant AI deployments, adaptation is concentrated at high software layers (model parameters and software-level state) while lower layers (instruction set, microarchitecture, hardware) are treated as fixed during deployment. [^2] Lower layers can carry mutable state (data, caches, firmware), and self-modifying substrates exist; the claim here is about the prevailing engineering pattern in mainstream ML systems. In contrast, biological multiscale competency architectures (MCA) delegate goal-directed control across scales, from cells to organs to organisms [34, 59]. I model this difference with Stack Theory, formalising systems as stacks of abstraction layers. Delegation then measures how far adaptive control reaches down the stack.
@@
The somewhat provocative claim this paper will make is that biological self-organising systems are more intelligent (in the efficiency-of-adaptation sense) than contemporary, predominantly fixed-stack implementations of AI. This is not a claim about all possible future AI. One aim of the paper is to identify design principles for building more adaptable hybrid and cyberphysical systems by shifting where adaptation occurs in the stack. This is because biological self-organisation delegates adaptation to lower levels of abstraction (for the generalisation of “levels of abstraction” beyond computing, see ESM Supplementary Note S2). Formally, I measure adaptability by the weakness of the constraints a system embodies [20, 16]. I then prove The Law of the Stack (Theorem 5.1), which is that utility (adaptability) achievable one layer up is bounded by the weakness of the realised policy one layer down. Finally, I link this to the FEP. Under a simple viability prior, maximising weakness is equivalent to minimising variational free energy (see ESM “Viability, lived history, and variational free energy”), which makes delegation a prerequisite for sustained free energy minimisation across scales.
@@
If these restrictions render the system over-constrained at $v$ (i.e. $\widehat{\Pi}_{\omega}=\emptyset$), then no common policy exists at that level of abstraction. If at least one part remains locally viable (i.e. $\widehat{\Pi}_{\alpha_{j}}\neq\emptyset$ for some $j$), then any continued function must proceed via a proper subset of parts whose restricted policy sets still admit a non-empty intersection. I refer to this formal situation as “cancer-like” purely by analogy.
@@
Hybrid agents (e.g. organoids, soft robots, human–AI organisations) are best treated as boundary-condition design problems, which I formalise using Stack Theory. One must engineer constraints that prune low-level policy spaces just enough to realise target collective behaviour while preserving a non-empty intersection of component policy sets (collective identity). Formalised this way, the optimal choice of boundary conditions are the weakest sufficient constraints to achieve the desired hybrid functionality.
@@ 2 Abstraction, Distribution and Delegation
In proposing Stack Theory [19], I extend this idea beyond computing. I illustrate abstraction, distribution, and delegation with limiting-case examples from computing, biology, and human organisations.
@@ 2.1 Abstraction
Abstraction layers are interfaces that hide lower-level details. In computation, Python is interpreted by C, which runs on assembly and machine code on hardware. Each higher-layer behaviour must be realised by some lower-layer behaviour, but the same abstract interface can be realised by multiple implementations. This many-to-one mapping is where time, energy and information costs hide, so the abstraction layer through which an agent senses and acts constrains its adaptability. In Stack Theory, useful invariants are treated as higher-level task constraints even when micro-implementations vary. Even hardware itself can be framed as an abstraction layer over the local environment [19].
Think of a stack as a directed graph of constraints. Higher-level behaviours are implemented by and limited by lower-level behaviours. In biology, cells enact organs, organs enact organisms, and organisms enact still more abstract behaviours such as communication and tool use [59, 16].
@@ 2.2 Distribution
Distribution refers to how work is spread within a layer. In many biological systems, lower levels of abstraction are literally smaller parts [32], so “going down” is like zooming in. Cells form organs and organisms form organisms. For human organisations, soldiers form squads and squads form platoons.
@@ 2.3 Delegated Control
Biological collectives differ in that cells and tissues are autonomously goal-directed at their own scale [6]. Higher-level coherence arises when the viable policy sets of parts intersect, producing a collective policy (collective identity). Control can propagate bottom-up and top-down by which I mean collectives constrain their parts (e.g. via signals or incentives), but parts also shape collective behaviour. For illustrative purposes, consider how members of a collective are constrained by the collective, for example members of a human society are incentivised to obey laws. This is a form of top-down control. Bottom-up then becomes just whatever the individual humans do within those constraints, even though the system as a whole might have emerged entirely bottom-up from anarchy to begin with. Each part of a collective acts on its own, and its actions imply a ‘policy’. Each individual policy determines how that part interacts with the other parts, or more to the point their policies. These interactions cause the system as a whole to act, which implies a collective policy. If collective and individual policies align in service of homeostasis, then the system can persist. It has a collective identity or goal. That identity can constrain parts of the collective. For example, I might say the level of abstraction is cells, and then describe an organ as a collective policy of those cells. Likewise, I might look at the level of organs, and describe a human as a collective policy of those organs. Or I could describe the human as the collective policy of cells, and see the organs as a sort of middle management layer.
@@
Economies illustrate useful limit cases to understand delegation. One might think of the people in the economy as parts at lower level of abstraction. Central planning concentrates control at the top, while extreme decentralisation delegates control to the bottom. In the central planning limit case, the population does work but has no control. In the limit those people could adapt given local information unavailable to the central planner [44], but central planner’s edicts constrain them. In the opposite extreme there is no coordination.
@@
Real systems occupy a continuum. Of course, there are limits on what sort of physical forms can store or process information in service of a goal [81], and this may limit how much I can distribute work or delegate control. Delegating control often coincides with shifting to finer granularity and greater distribution. However, as computers illustrate, abstraction can also change without changing the spatial scale of components (Figures 1 and 2).
@@
Figure 1: Control or agency and thus adaptation in a given system can take place at a particular level of abstraction. For example, I can program in Python, or I can go down a level and program in C. If I program in C it doesn’t matter whether I program one computer or fifty. Either way my control is at the same level of abstraction (figure from [19]).
@@
Figure 2: Work, resources and so on can be distributed across a system. Distribution is distinct from delegation of control. For example, both food stamps and universal basic income might distribute the same amount of resources to the same people within an economic system, but the former more tightly constrains the recipients in how they use those resources (figure from [19]).
@@ 2.4 Reasons to Suspect Delegation Facilitates Adaptation
Theorem 5.1 provides the formal claim. Here I briefly motivate why delegation should matter. Because perceived complexity is defined by the abstraction layer through which a system interacts with its environment [17], measures of intelligence that ignore embodiment are incomplete; an “objective” measure must be enactive [86].
First, adaptation closer to the substrate can exploit more degrees of freedom. Cutting out an abstraction layer can yield large efficiency gains (C vs Python) [65]; reconfigurable substrates such as FPGAs push adaptation further down the stack [63].
Second, delegation can improve robustness via graceful degradation, avoiding single points of failure [55]. For example, to control a drone swarm I could single out one drone to be the ‘leader’, and have it run a program that co-ordinates the other drones. However that would make the swarm vulnerable. If the leader drone were damaged, the swarm would de-cohere. In contrast if I can get the swarm to co-ordinate through entirely decentralised control, individual drones negotiating with those around them, then the swarm can remain coherent regardless of how many drones are disabled. This sort of decentralised swarm control can also be observed in nature, such as with flocks of starlings which fly in complex formations [25]. In other words, the swarm ‘adapts’ to adverse conditions.
Third, in human organisations delegation shortens feedback loops. In military doctrine, “mission command” delegates decision-making to units closest to the situation, granting latitude to deviate from detailed orders in service of higher-level intent [79]. This trades central control for responsiveness. See ESM Note S7 for a concrete historical illustration (Nordbat 2 [47, 48]).
@@ 3 What is Adaptation?
To make accurate inferences a system must be able to tell the difference between what it has caused, and what it has merely observed [60, 72, 31]. It is often difficult to determine cause. ‘Confounding’ occurs when an event appears to be the cause of the effect we’re trying to predict, when in truth it is merely correlated. This can lead one to draw erroneous conclusions.
@@
By actively intervening in the environment, a system can experiment to establish cause [^3] Assuming such experimentation is enough to mitigate ambiguity and eliminate unknowns, and that a heuristic such as weakness is applied to ensure the system converges on variable definitions which represent cause accurately [14].. Then, it can construct a causal graph. There are two ways to do that. One option would be to assume the nodes (variables, actions and other abstract objects and properties), and then learn the causal relations between them. This might work well in theory [74], but it requires foreknowledge of objects and properties between which causal relations can be established. The alternative is to assume the causal relations (the edges in the graph), and then learn the variables [14, 19, 18]. This needs only attraction and repulsion from physical states [12, 19], otherwise known as valence. I take this latter approach. While causal discovery of this sort remains difficult to engineer in open-ended settings (e.g. unobserved confounders and combinatorial hypothesis spaces), I treat this as a modelling stance and focus on the stack-level delegation constraint formalised in Theorem 5.1.
@@
Learning nodes to fit a graph is like figuring out which computer program was used to generate a sequence. Assume I have fed a set of inputs into a computer. For each input the computer responded with a particular output. The computer runs a ‘policy’ program which maps inputs to outputs [^4] I call such programs policies, but this should not be confused with the more specific definition policy in reinforcement learning [84].. That policy is the cause of the input-output history observed so far. I want to know the policy that caused historical outputs so that I can use it to predict future outputs. The problem is that more than one policy could have produced the history we’ve observed, and each of those policies produce different outputs when I give the computer inputs it hasn’t encountered before. From a naive Bayesian perspective, all of them are equally likely to have caused the history I saw. If I wish to know which policy actually caused the outputs, I need a way to whittle down the remaining possibilities.
@@
There are two ways to compare policies that look the same with respect to historical data, but differ with respect to possible future data [17]. One is to compare their form, and choose policies which are simpler in the sense of being more compressible. However, in an interactive setting simplicity is an unreliable indicator of plausibility [56, 17]. The alternative is to compare policies by their function, in terms of how they behave given inputs for which I do not have historical data. Adaptive systems must satisfy functional constraints [83]. I can compare systems by these constraints. Each policy implies a constraint on possible outputs given a set of possible inputs. The weaker that constraint is, the more variety of future behaviour it permits. Put another way, if two policies satisfy a goal, then the weaker policy satisfies it in more possible worlds [^5] To the extent that such possible worlds can be differentiated from one another within the subjective constraint of an abstraction layer.. Previous research showed that if two policies are identical with respect to historical data, then the policy most likely to have generated historical data is the weaker [21, 19] (see ESM Supplementary Note S3 for formal details.) of the two [14, 20, 18]. The policy most likely to have caused behaviour is the weakest [14], and in experiments a system which preferred weaker policies outperformed one that preferred the simpler policies by $110-500\%$ [15]. This has been used to establish an upper bound on embodied ‘intelligence’ [16]. I will use the term ‘w-maxing’ to mean a system optimises for weak policies, and ‘simp-maxing’ to mean a system optimises for simpler (minimum description length [75]) policies, to conform to the terminology used in previous work [22, 19, 21]. Put another way, if two hypotheses appear equally good with respect to past data, the w-maxing system will use policy weakness as a heuristic to decide which policy to use, whereas a simp-maxing system will use simplicity.
@@
Weakness is not the same thing as simplicity. Simplicity is a property of a policy’s description in a chosen encoding. Weakness is a property of the policy’s remaining completion space inside the embodied language $L_{\mathfrak{v}}$. Two correct policies can therefore have very different weakness even when their descriptions are equally short, or when one is syntactically longer. Weakness is also not mere truth set size. It counts how many distinct further commitments remain available to the agent given what it can express and implement at that abstraction layer. In this sense, w-maxing implements a least commitment heuristic stated at the level of function, rather than at the level of syntax. Since the completion structure depends on the vocabulary, weakness naturally incorporates embodiment. Changing sensors, effectors, or internal degrees of freedom changes $\mathfrak{v}$ and changes which refinements count as distinct. In the uniform case where all remaining completions are treated as equally plausible, maximising weakness coincides with a maximum entropy preference over viable futures.
@@
An abstraction layer can be understood as a border between form and function. For example a program takes a form when it is written in a language, but when it is run on a computer it has a particular function or effect upon the world. Simp-maxing is a matter of form, while w-maxing is a matter of function [22]. When we account for interpretation between form and function it becomes clear that software is a physical thing, like a tool. A tool can serve more than one purpose. For example, a hammer is both a weapon and a paperweight. This is what “weakness” measures. A tool that implies a weaker constraint can complete more tasks [10] than another tool. By preferring weaker policies, one avoids unnecessary commitments, meaning the weakest policies complete a ‘wider range of tasks’ than others (see ESM Note S3 for discussion of weakness and related connections (e.g. maximum entropy).). To complete a ‘wider range of tasks’ in the embodied sense, a tool must be both more sample and more energy efficient, and so the weakest policy is the most efficient in both senses. Put another way, the same physical matter can simultaneously contribute to many different tasks or computations superimposed over each other at different scales. A ‘weaker’ policy serves more masters.
@@
ML systems do not typically optimise for either weaker constraints on function (w-maxing), or simplicity of form (simp-maxing). This may explain some of their relative inefficiency [13, 15]. Even if they were w-maxing, the capacity for w-maxing is severely limited by the abstraction layer. An abstraction layer can be a “causal isolator” [35] between a software “mind” at higher levels of abstraction, and its environment at lower levels. Embodiment by another name. Different abstraction layers permit different upper bounds on policy weakness, in the same way that two bodies might differ in how able they are to achieve a given goal. Different embodiments give different upper bounds on ‘intelligence’ [16].
@@
That biological self-organising systems adapt efficiently means they are able to embody weaker policies than today’s ML systems. This capacity hinges on the abstraction layer [16]. An abstraction layer is like a window through which only part of the world can be seen. It conveys an inductive bias that helps or hinders adaptation. When I optimise the system at a lower level of abstraction, this is like saying I can “move” this window to convey a different inductive bias. The bigger the window, the “weaker” the policy it is possible to learn. What this means is that intelligence is primarily a matter of the lens through which we interact with a problem. Hardware, not just software. In enactivist terms, an intelligent choice of abstraction layer is one that reduces the “big world” of all information to a “small world” of relevant information [95, 94, 93]. In a typical computer, lower levels of abstraction tend to be static. The window cannot move. I will show a static abstraction layer limits adaptability at higher levels of abstraction by limiting how weak the policies embodied by the system can be. Adaptability at high levels requires adaptability at low levels, and delegation of control facilitates this.

### v7

# Hybrid Agency, Artificial Intelligence, Cyberphysical Systems

@@ Abstract
Are biological self-organising systems more “intelligent” than artificial intelligence (AI)? If so, why? I address this question using a mathematical framework that defines intelligence in terms of adaptability. Systems are modelled as stacks of abstraction layers (Stack Theory) and compared by how effectively they delegate agentic control down their stacks. I illustrate this using computational, biological, military, and economic systems. Contemporary AI typically relies on static, human-engineered stacks whose lower layers are fixed during deployment. Put provocatively, such systems resemble inflexible bureaucracies that adapt only top-down. Biological systems are more intelligent because they delegate adaptation. Formally, I prove three results. The first theorem is The Law of the Stack, which shows adaptability at higher layers is bottlenecked by adaptability at lower layers. Second, I show maximising adaptability is equivalent to minimising variational free energy under viability assumptions, proving delegation is necessary for free-energy minimisation. Finally, I generalise bioelectric accounts of cancer to show cancer-like failures can occur in non-biological systems when delegation is inadequate. These yield design principles for building robust systems via delegated control, re-framing hybrid agents (e.g. organoids or human–AI systems) as weak boundary-condition design problems in which constraints shape low-level policy spaces while preserving collective identity.
keywords:
Stack Theory, delegation, w-maxing, adaptability, hybrid agency
@@ 1 Introduction
Organoids are living systems whose behaviour is sculpted by engineered constraints. Human brain cells embedded in a computational interface have been trained to play Pong, achieving remarkable sample efficiency compared to a contemporary reinforcement learning agent [khajehnejad2024]. Similarly, Physarum slime mould is normally an indiscriminate expander, but it reliably computes shortest paths when placed inside a maze [nakagaki2000]. These behaviours are neither programmed nor accidental. They arise because biological agency is coupled to carefully designed boundary conditions. Such systems are hybrid agents, jointly produced by bottom-up cellular dynamics and top-down environmental structure [sole2019].
@@
Biological systems self-organise across scales. Their parts and environment interact to produce spatiotemporal order [ashby1947, seeley2002, rosas2018, bell2007, kelso1997, Friston2010, tognoli2014]. Many systems proactively learn, storing information about the past to anticipate future needs [kukushkin2024]. To adapt in complex environments, a system must infer cause and effect [pearl2018, merker2005, barron2016]. Throughout, I use ‘intelligence’ in the specific sense of the sample and energy efficiency of adaptation [bennett2024a, chollet2019, bennettmaruyama2022b, bennett2023c] [^1] Mainstream AI typically treats intelligence as performance across tasks. My conclusions concern efficiency of acquiring said performance. Accordingly, claims about “AI” should be read as applying to dominant contemporary ML training and deployment stacks, rather than all possible systems. See ESM Note S1..
@@
Two common explanations for biological efficiency are active inference under the Free Energy Principle (FEP) [Friston2010, friston2013, parr2024] and powerful inductive biases [chollet2019]. These address, respectively, how agents update beliefs and act, and what structure is pre-specified. This paper isolates a complementary structural question, namely where in a system is adaptation permitted?
@@
Computers are organised into layers of abstraction. In dominant AI deployments, adaptation is concentrated at high software layers (model parameters and software-level state) while lower layers (instruction set, microarchitecture, hardware) are treated as fixed during deployment. [^2] Lower layers can carry mutable state (data, caches, firmware), and self-modifying substrates exist. The claim only pertains to the prevailing engineering pattern in mainstream ML systems. In contrast, biological multiscale competency architectures (MCA) delegate goal-directed control across scales, from cells to organs to organisms [fields2020, levin2024]. I model this difference with Stack Theory, formalising systems as stacks of abstraction layers. Delegation then measures how far adaptive control reaches down the stack.
@@
The somewhat provocative claim this paper will make is that biological self-organising systems are more intelligent (in the efficiency-of-adaptation sense) than contemporary, predominantly fixed-stack implementations of AI. This is not a claim about all possible future AI. Delegation is also not a binary distinction between “static AI” and “dynamic biology”. It is a continuum. Many deployed AI systems concentrate adaptation into an offline training phase, whereas biological systems distribute plasticity and regulation across morphology, metabolism and development. One aim of the paper is to identify design principles for building more adaptable hybrid and cyberphysical systems by shifting where adaptation occurs in the stack. This is because biological self-organisation delegates adaptation to lower levels of abstraction (for the generalisation of “levels of abstraction” beyond computing, see ESM Supplementary Note S2). Formally, I measure adaptability by the weakness of the constraints a system embodies [bennett2025d, bennett2024a]. I then prove The Law of the Stack (Theorem 5.1), which is that utility (adaptability) achievable one layer up is bounded by the weakness of the realised policy one layer down. Finally, I link this to the FEP. Under a simple viability prior, maximising weakness is equivalent to minimising variational free energy (see ESM “Viability, lived history, and variational free energy”), which makes delegation a prerequisite for sustained free energy minimisation across scales.
@@
If these restrictions render the system over-constrained at $v$ (i.e. $\widehat{\Pi}_{\omega}=\emptyset$), then no common policy exists at that level of abstraction. If at least one part remains locally viable (i.e. $\widehat{\Pi}_{\alpha_{j}}\neq\emptyset$ for some $j$), then any continued function must proceed via a proper subset of parts whose restricted policy sets still admit a non-empty intersection. I refer to this formal situation as “cancer-like” purely by analogy (see Definition 6.1 and Proposition 6.1).
@@
Hybrid agents (e.g. organoids, soft robots, human–AI organisations) are best treated as boundary-condition design problems, which I formalise using Stack Theory [^3] This paper’s use of “Stack Theory” is a formal model of layered control and delegation, distinct from and complementary to Bratton’s geopolitical use of “The Stack” [bratton2016].. One must engineer constraints that prune low-level policy spaces just enough to realise target collective behaviour while preserving a non-empty intersection of component policy sets (collective identity). Formalised this way, the optimal choice of boundary conditions is the weakest set of sufficient constraints to achieve the desired hybrid functionality.
@@ 2 Abstraction, Distribution and Delegation
In proposing Stack Theory [bennett2025thesis], I extend this idea beyond computing. I illustrate abstraction, distribution, and delegation with limiting-case examples from computing, biology, and human organisations.
@@ 2.1 Abstraction
Abstraction layers are interfaces that hide lower-level details. In computation, Python is interpreted by C, which runs on assembly and machine code on hardware. Each higher-layer behaviour must be realised by some lower-layer behaviour, but the same abstract interface can be realised by multiple implementations. This many-to-one mapping is where time, energy and information costs hide, so the abstraction layer through which an agent senses and acts constrains its adaptability. In Stack Theory, useful invariants are treated as higher-level task constraints even when micro-implementations vary. Even hardware itself can be framed as an abstraction layer over the local environment [bennett2025thesis].
Think of a stack as a directed graph of constraints. Higher-level behaviours are implemented by and limited by lower-level behaviours. In biology, cells enact organs, organs enact organisms, and organisms enact still more abstract behaviours such as communication and tool use [levin2024, bennett2024a].
@@ 2.2 Distribution
Distribution refers to how work is spread within a layer. In many biological systems, lower levels of abstraction are literally smaller parts [delgado1997], so “going down” is like zooming in. Cells form organs and organisms form organisms. For human organisations, soldiers form squads and squads form platoons.
@@ 2.3 Delegated Control
Biological collectives differ in that cells and tissues are autonomously goal-directed at their own scale [ball2023]. Higher-level coherence arises when the viable policy sets of parts intersect, producing a collective policy (collective identity). Control can propagate bottom-up and top-down by which I mean collectives constrain their parts (e.g. via signals or incentives), but parts also shape collective behaviour. For illustrative purposes, consider how members of a collective are constrained by the collective, for example members of a human society are incentivised to obey laws. This is a form of top-down control. Bottom-up then becomes just whatever the individual humans do within those constraints. Each part of a collective acts on its own, and its actions imply a ‘policy’. Each individual policy determines how that part interacts with the other parts, or more to the point their policies. These interactions cause the system as a whole to act, which implies a collective policy. If collective and individual policies align in service of homeostasis, then the system can persist.
@@
Economies illustrate useful limit cases to understand delegation. One might think of the people in the economy as parts at lower level of abstraction. Central planning concentrates control at the top, while extreme decentralisation delegates control to the bottom. In the central planning limit case, the population does work but has no control. In the limit those people could adapt given local information unavailable to the central planner [hayek1945use], but central planner’s edicts constrain them. In the opposite extreme there is no coordination.
@@
Real systems occupy a continuum. Of course, there are limits on what sort of physical forms can store or process information in service of a goal [sole2024], and this may limit how much I can distribute work or delegate control. Delegating control often coincides with shifting to finer granularity and greater distribution. However, as computers illustrate, abstraction can also change without changing the spatial scale of components (Figures 1 and 2).
@@
Figure 1: Control or agency and thus adaptation in a given system can take place at a particular level of abstraction. For example, I can program in Python, or I can go down a level and program in C. If I program in C it doesn’t matter whether I program one computer or fifty. Either way my control is at the same level of abstraction (figure from [bennett2025thesis]).
@@
Figure 2: Work, resources and so on can be distributed across a system. Distribution is distinct from delegation of control. For example, both food stamps and universal basic income might distribute the same amount of resources to the same people within an economic system, but the former more tightly constrains the recipients in how they use those resources (figure from [bennett2025thesis]).
@@ 2.4 Reasons to Suspect Delegation Facilitates Adaptation
Theorem 5.1 provides the formal claim. Here I briefly motivate why delegation should matter. Because perceived complexity is defined by the abstraction layer through which a system interacts with its environment [bennett2024b], measures of intelligence that ignore embodiment are incomplete. An “objective” measure must be enactive [thompson2007].
First, adaptation closer to the substrate can exploit more degrees of freedom and specialisation. Programming at a lower level of abstraction can yield large efficiency gains (e.g. C vs Python) [oden2020]. Reconfigurable substrates such as FPGAs can push adaptation further down the stack, enabling even more specialisation and efficiency [nguyen2020].
Second, delegation can improve robustness via graceful degradation, avoiding single points of failure [kuniavsky2010]. For example, to co-ordinate a drone swarm I could program one drone to plan the movements of the others. However, if that leader drone were damaged, the swarm would fail to co-ordinate. A more resilient approach is to program the swarm in a decentralised manner where each drone negotiates solutions with its neighbours, so the swarm has no single point of failure. Decentralised swarm control also emerges naturally, such as with flocks of starlings flying in complex formations [cavagna2010].
Third, in human organisations delegation shortens feedback loops. In military doctrine, “mission command” delegates decision-making to units closest to the situation, granting latitude to deviate from detailed orders in service of higher-level intent [Shamir01102010]. This trades central control for responsiveness. See ESM Note S7 for a concrete historical illustration (Nordbat 2 [ingesson2016, ingesson2017]).
@@ 3 What is Adaptation?
To make accurate inferences a system must be able to tell the difference between what it has caused, and what it has merely observed [merker2005, pearl2018, dawid2002]. It is often difficult to determine cause. ‘Confounding’ occurs when an event appears to be the cause of the effect we’re trying to predict, when in truth it is merely correlated. This can lead one to draw erroneous conclusions.
@@
By actively intervening in the environment, a system can experiment to establish cause [^4] Assuming such experimentation is enough to mitigate ambiguity and eliminate unknowns, and that a heuristic such as weakness is applied to ensure the system converges on variable definitions which represent cause accurately [bennett2023c].. Then, it can construct a causal graph. There are two ways to do that. One option would be to assume the nodes (variables, actions and other abstract objects and properties), and then learn the causal relations between them. This might work well in theory [richens2024robust], but it requires foreknowledge of objects and properties between which causal relations can be established. The alternative is to assume the causal relations (the edges in the graph), and then learn the variables [bennett2023c, bennett2025thesis, bennett2025e]. This needs only attraction and repulsion from physical states [bennettwelshciaunica2024, bennett2025thesis], otherwise known as valence. I take this latter approach. While causal discovery of this sort remains difficult to engineer in open-ended settings (e.g. unobserved confounders and combinatorial hypothesis spaces), I treat this as a modelling stance and focus on the stack-level delegation constraint formalised in Theorem 5.1.
@@
Learning nodes to fit a graph is like figuring out which computer program was used to generate a sequence. Assume I have fed a set of inputs into a computer. For each input the computer responded with a particular output. The computer runs a ‘policy’ program which maps inputs to outputs [^5] I call such programs policies, but this should not be confused with the more specific definition policy in reinforcement learning [sutton2018].. That policy is the cause of the input-output history observed so far. I want to know the policy that caused historical outputs so that I can use it to predict future outputs. The problem is that more than one policy could have produced the history we’ve observed, and each of those policies produce different outputs when I give the computer inputs it hasn’t encountered before. From a naive Bayesian perspective, all of them are equally likely to have caused the history I saw. If I wish to know which policy actually caused the outputs, I need a way to whittle down the remaining possibilities.
@@
There are two ways to compare policies that look the same with respect to historical data, but differ with respect to possible future data [bennett2024b]. One is to compare their form, and choose policies which are simpler in the sense of being more compressible. However, in an interactive setting simplicity is an unreliable indicator of plausibility [leike2015, bennett2024b]. The alternative is to compare policies by their function, in terms of how they behave given inputs for which I do not have historical data. Adaptive systems must satisfy functional constraints [sole2022]. I can compare systems by these constraints. Each policy implies a constraint on possible outputs given a set of possible inputs. The weaker that constraint is, the more variety of future behaviour it permits. Put another way, if two policies satisfy a goal, then the weaker policy satisfies it in more possible worlds [^6] To the extent that such possible worlds can be differentiated from one another within the subjective constraint of an abstraction layer.. Previous research showed that if two policies are identical with respect to historical data, then the policy most likely to have generated historical data is the weaker [bennett2024appendices, bennett2025thesis] (see ESM Supplementary Note S3 for formal details.) of the two [bennett2023c, bennett2025d, bennett2025e]. The policy most likely to have caused behaviour is the weakest [bennett2023c], and in experiments a system which preferred weaker policies outperformed one that preferred the simpler policies by $110-500\%$ [bennett2023b]. This has been used to establish an upper bound on embodied ‘intelligence’ [bennett2024a]. I will use the term ‘w-maxing’ to mean a system optimises for weak policies, and ‘simp-maxing’ to mean a system optimises for simpler (minimum description length [rissanen1978]) policies, to conform to the terminology used in previous work [bennett2025b, bennett2025thesis, bennett2024appendices]. Put another way, if two hypotheses appear equally good with respect to past data, the w-maxing system will use policy weakness as a heuristic to decide which policy to use, whereas a simp-maxing system will use simplicity.
@@
Weakness is not the same thing as simplicity. Simplicity is a property of a policy’s description in a chosen encoding. The weakness of a policy is the degrees of freedom remaining to a system that embodies that policy. Two correct policies can therefore have very different weakness even when their descriptions are equally short, or when one is syntactically longer. Weakness is also not mere truth set size. It counts how many distinct further commitments remain available to the agent given what it can express and implement at that abstraction layer. In this sense, w-maxing implements a least commitment heuristic stated at the level of function, rather than at the level of syntax. Since the completion structure depends on the vocabulary, weakness naturally incorporates embodiment. Changing sensors, effectors, or internal degrees of freedom changes $\mathfrak{v}$ and changes which refinements count as distinct. In the uniform case where all remaining completions are treated as equally plausible, maximising weakness coincides with a maximum entropy preference over viable futures.
@@
An abstraction layer can be understood as a border between form and function. For example a program takes a form when it is written in a language, but when it is run on a computer it has a particular function or effect upon the world. Simp-maxing is a matter of form, while w-maxing is a matter of function [bennett2025b]. When we account for interpretation between form and function it becomes clear that software is a physical thing, like a tool. A tool can serve more than one purpose. For example, a hammer is both a weapon and a paperweight. This is what “weakness” measures. A tool that implies a weaker constraint can complete more tasks [bennettmaruyama2022a] than another tool. By preferring weaker policies, one avoids unnecessary commitments, meaning the weakest policies complete a ‘wider range of tasks’ than others (see ESM Note S3 for discussion of weakness and related concepts, such as maximum entropy). To complete a ‘wider range of tasks’ in the embodied sense, a tool must be both more sample and more energy efficient, and so the weakest policy is the most efficient in both senses. Put another way, the same physical matter can simultaneously contribute to many different tasks or computations superimposed over each other at different scales. A ‘weaker’ policy serves more masters.
@@
ML systems do not typically optimise for either weaker constraints on function (w-maxing), or simplicity of form (simp-maxing). This may explain some of their relative inefficiency [bennett2022a, bennett2023b]. Even if they were w-maxing, the capacity for w-maxing is severely limited by the abstraction layer. An abstraction layer can be a “causal isolator” [friston2023] between a software “mind” at higher levels of abstraction, and its environment at lower levels. Embodiment by another name. Different abstraction layers permit different upper bounds on policy weakness, in the same way that two bodies might differ in how able they are to achieve a given goal. Different embodiments give different upper bounds on ‘intelligence’ [bennett2024a].
@@
That biological self-organising systems adapt efficiently means they are able to embody weaker policies than today’s ML systems. This capacity hinges on the abstraction layer [bennett2024a]. An abstraction layer is like a window through which only part of the world can be seen. It conveys an inductive bias that helps or hinders adaptation. When I optimise the system at a lower level of abstraction, this is like saying I can “move” this window to convey a different inductive bias. The bigger the window, the “weaker” the policy it is possible to learn. What this means is that intelligence is primarily a matter of the lens through which we interact with a problem. Hardware, not just software. In enactivist terms, an intelligent choice of abstraction layer is one that reduces the “big world” of all information to a “small world” of relevant information [vervaeke2012, vervaeke2013a, vervaeke2013b]. In a typical computer, lower levels of abstraction tend to be static. The window cannot move. I will show a static abstraction layer limits adaptability at higher levels of abstraction by limiting how weak the policies embodied by the system can be. Adaptability at high levels requires adaptability at low levels, and delegation of control facilitates this.